OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Calculate the geometric mean of 3 and 9.

OpenStudy (solomonzelman):

So, a1=3 and a3=9 and you are missing a2. \(\large\color{black}{ a_1\times r^{3-1}=a_3}\) \(\large\color{black}{ 3\times r^{2}=9}\) Solve for r.

OpenStudy (solomonzelman):

I really want to say though, that you will have to answers for a2 (the messing geometric mean)

OpenStudy (anonymous):

sqrt(3*9)=3root 3

OpenStudy (solomonzelman):

excuse my grammar

OpenStudy (perl):



OpenStudy (solomonzelman):